Andrew McAward and Ed Schwendemann – Final Paper

Your presentation to the class was very good, and you had some good background information on the Blackstone River that other groups did not mention.

The final project paper is also very good. You did a reasonable job with the measurements of the different categories of surfaces, and the discussion of you would define and compute the run-off index is also mostly good. The main comment I have is that there are a number of other ways to set up reasonable methods there and some of them might be better choices in some ways. For instance, the way you computed z-scores for each block in each category, then took a weighted average of the z-scores is actually a bit “backwards” if you want to use statistical techniques based on the standard normal distribution to identify the problem areas. Namely, it would be better to take the weighted average of the areas of the the different categories of surfaces for each “block” and call that the RI for that block. Then if you computed the average and SD of those RI values and did the z-scores based on the single RI values, you would have a direct basis of comparison for the different blocks. The way you set this up would not yield values that followed a standard normal distribution (even approximately). So you would still need to do something else to compare your runoff-index values for different blocks and pick out the problem areas. You did not really consider that last step, so what you said was not quite complete.

The other general comment I have is that the paper has two very abrupt transitions, one at the point where you begin to describe the campus measurement process, and the second at the start of the section about ideas for things that could be done on campus to minimize runoff problems. The “seams are showing” in both of those places. You could have either added section headings to indicate where you were going next, or included transition sentences to do the same thing.

Technical points in the citations and bibliography are handled very well. Specific Comments

1) “Decrease in slope” is not correct here. You mean “change in elevation.” The slope is the ratio “rise over run” and that is actually pretty consistent over the whole hillside.

2) The quotation about the Blackstone River being “the most polluted river in the country with respect to toxic sediments” is actually used twice in three pages. This is repetitive and the second does not add anything. It might have been good to mention exactly why this is important in combination with runoff, though. Note that water flowing into the Blackstone can stir up sediment from the river bed and mix it back into the water.

3) “Activities like boating” is not a complete sentence. You could say “ ... suitable for ‘secondary contact’ activities only” (such as boating).

4) This is one of the “seams are showing” spots I mentioned above.

5) I’m assuming the w_{i} are the weighting factors discussed two pages before. It might be good to mention that that is what they are before giving the formula – say “Let w_{i} be the weighting factors introduced previously. Then we compute a runoff index value by ... ”

6) This is the other “seams are showing” place.

Final Project Grades Presentation: A- (92) Paper Content: A- (92) Paper Form: A- (92)